4. Find the ratio in which the line segment joining the points (-3, 10) and (6, – 8) is divided by (-1, 6).

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Consider the ratio in which the line segment joining ( -3, 10) and (6, -8) is divided by point ( -1, 6) be k :1. Therefore, -1 = ( 6k-3)/(k+1) –k – 1 = 6k -3 7k = 2 k = 2/7 Therefore, the required ratio is 2: 7.